OpenStudy (anonymous):

what is the vertex form of the equation? y = -x^2 + 12x - 4 and show me how you did it?

OpenStudy (anonymous):

@Ria23 can you help?

OpenStudy (anonymous):

@jordanloveangel ?

OpenStudy (anonymous):

@Loser66 or @ganeshie8

OpenStudy (dan815):

(y-k)=a(x-j)^2

OpenStudy (dan815):

thats vertex form

OpenStudy (anonymous):

Yeah I know. do you know how to convert it?

OpenStudy (dan815):

so complete the square..

OpenStudy (anonymous):

ugh Everyone always says that but i dont really know how to do that

OpenStudy (dan815):

y = -x^2 + 12x - 4 -y=x^2-12x+4 =(x-6)^2+4-36 -y +32 = (x-6)^2 (y-32)=-(x-6)^2

OpenStudy (dan815):

completing the square is

OpenStudy (dan815):

u add some number to both sides so that u can factor the x part into that square form

OpenStudy (dan815):